OpenStudy (anonymous):

find the sum of the first eight terms of the AP -1,2,5,8

OpenStudy (anonymous):

-1+2+5+8+(8+3)+(8+3+3)+(8+3+3+3)+(8+3+3+3+3)=SOMETHING USE CALCULATORZZZ

OpenStudy (cwrw238):

sum of n terms of an AP is given by Sn = (n/2)[2a + (n-1)d] where n = no. of terms, a = first term and d = common difference

OpenStudy (anonymous):

Or you could just add all those 3's

OpenStudy (cwrw238):

yes that would give the correct answer its ok for a small number of terms but heavy work say for 502 terms - then you use the formula!